OpenStudy (anonymous):

On a cloudy day, Yong Sun needed to know the height of a window in a building. Yong Sun positioned a mirror on the ground between herself and the building so that when he looked in the mirror, she saw the window. If the mirror was 22.88 cm from her feet and 34.15 m from the base of the building and Yong Sun's eye was 1.89 m above the ground, how high up on the building was the window located? Round to the nearest meter. A. 1 m B. 3 m C. 127 m D. 282 m

OpenStudy (anonymous):

similar triangles 22.88cm/1.89m=34.15m/x .2288m/1.89m=34.14m/x x=(34.14*1.89)/0.2288=282m